OpenStudy (dallaslifebaby01):

Translate the sentence into a formula. In a right triangle, the square of the measures of the hypotenuse c is equal to the sum of the squares of the measures of the legs, a and b

OpenStudy (welshfella):



OpenStudy (dallaslifebaby01):

i need it in a formula @welshfella ( cool user name btw )

OpenStudy (welshfella):

c is the hypotenuse and a and b are the legs so c^2 = what?

OpenStudy (dallaslifebaby01):

I'm not sure what that would equal... would it be c^2=AB?

OpenStudy (dallaslifebaby01):

@welshfella

OpenStudy (welshfella):

No each of the legs are squared ( c^2 = 'c squared' and the squares of a and b are added

OpenStudy (dallaslifebaby01):

So it would be C^2 = A^2 + B^2 ? sorry if its not and I'm not getting it... I'm bad at maths

OpenStudy (welshfella):

Thats it - although sides are usually written in lower case c^2 = a^2 + b^2

OpenStudy (dallaslifebaby01):

oh okay thank you for your help!

OpenStudy (welshfella):

angles are usually written in uppercase. welcome

OpenStudy (mathmale):

Please look up "Pythagorean Theorem." How is that related to your question?